OpenStudy (anonymous):

What would -1/2 /8 be???

OpenStudy (anonymous):

-1/2 / 8 = -1/2 *1/8= -1/16

OpenStudy (anonymous):

okay, thanks so much!

OpenStudy (anonymous):

Do note that (-1/2)/8 is different from -1/(2/8)

OpenStudy (anonymous):

thx for the tip(: